OpenStudy (anonymous):

2. In this problem, you will investigate the family of function h(x)= x+cos(ax) where a is a positive constant such that 0 < a < 4 A. For what value of a will h(x) have a relative maximum at x=1? B. For what value of a will h(x) have an inflection point at x=1? C. Which values of a make h(x) strictly increasing? Justify your answer. Another non-test but need to know how to solve question.

OpenStudy (anonymous):

I will definately give medal for this question. I need to know how to solve these questions.

OpenStudy (anonymous):

the problem is that i have no clue at all how to solve such a question.... never seen it before.

OpenStudy (anonymous):

@amistre64 Would it be okay if you help me with one more question?

OpenStudy (amistre64):

derivatives help; just treat "a" as a constant ... name it 5 if need be

OpenStudy (amistre64):

or 3 since its nags us to be between 0 and 4 :)

OpenStudy (anonymous):

so do the derivative of that formula? would the derivative be1-sin(ax)(a)?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

So then just plug in 1 and that would be the answer for 2. A right?

OpenStudy (amistre64):

now we want to determine of there is any place where the slope is zero, or undefined for this derivative; this gives us "critical points" to test with

OpenStudy (anonymous):

Okay... But I am not quite sure what I would need to be doing after this step.

OpenStudy (amistre64):

1 - a sin(ax) = 0 a sin(ax) = 1 sin(ax) = 1/a ax = arcsin(1/a) x = (arcsin(1/a))/a

OpenStudy (amistre64):

what is the largest "a" between 0 and 4 such that x=1?

OpenStudy (anonymous):

woo.. tough question.

OpenStudy (amistre64):

or rather what value of a produce and x=1

OpenStudy (amistre64):

x = (arcsin(1/a))/a 1 = (arcsin(1/a))/a a = arcsin(1/a) sin(a) = 1/a

OpenStudy (amistre64):

im thinking that maybe this function in "a" can be derived as well, but its an iffy thought

OpenStudy (amistre64):

a cos(a) + sin(a) = 0

OpenStudy (anonymous):

hum.... with the equation you gave me, I get this vague answer that say sin(a) = 1/a right? ... Oh then sin(a)= 1/a so then cos(a) would basically mean -1/a right?

OpenStudy (amistre64):

let me rethink this; since the value of a depends on x (x=1) to be exact then we can consider a implicitly as a function of x \[x+cos(ax)\] \[1-sin(ax)(a'x+a)=0~:~x=1\] \[1-sin(ax)(a'+a)=0\] and do the algebra shuffle

OpenStudy (amistre64):

\[1=sin(a)(a'+a)\] \[csc(a)=a'+a\] \[a'=csc(a)-a\]

OpenStudy (amistre64):

which i believe is the same thing as sin(a) = 1/a

OpenStudy (amistre64):

im at a loss on it at the moment ....

OpenStudy (anonymous):

I can wait haha take your time. ty

OpenStudy (amistre64):

i was thinking that the interval on x was from 0 to 4, we only need to determine a suitable value of a such that the slope of the line at x=1 is 0, relative to any epsilon beside it

OpenStudy (amistre64):

what level is this?

OpenStudy (amistre64):

h(x)= x+cos(ax) , assume a to be constant h' = 1 - a sin(ax) = 0, at x=1 0 = 1 - a sin(a) 1/a = sin(a) i get 2 as by graphing that fit, one is sure to be a min and the other a max

OpenStudy (anonymous):

haha this is just calculus ab class.

OpenStudy (anonymous):

so do we do the second derivative test now for this?

OpenStudy (anonymous):

I think 1.114 would be the max because the original curve has highest point here?

OpenStudy (amistre64):

yeah, we can try that, it might actually help us out ... might not :) h' = 1 - a sin(ax) h'' = a^2 cos(ax)

OpenStudy (amistre64):

yes, but how to determine an exact value for a instead of an approximation ....

OpenStudy (amistre64):

h'' = -a^2 cos(ax) that is

OpenStudy (anonymous):

I really appreciate your effort to really help me out.

OpenStudy (amistre64):

at x=1 thats h'' = -a^2 cos(a) we would need the property such that any value a little bit to the left and right of a is ?

OpenStudy (amistre64):

youre welcome so far :)

OpenStudy (anonymous):

yes. I think that when the left side is sloping upward and right side is sloping downward that would be the max?

OpenStudy (amistre64):

slope relates to h' how does concavity help us? the rate of change of the slope

OpenStudy (amistre64):

i always have to refer to y=x^2; since i know its cave up, and min y = x^2 y' = 2x y'' = 2, is positive ; so when y'' is negative, we have cave down and max

OpenStudy (amistre64):

since -a^2 is always negative, we would want to confine cos(a) such that: 0 <= a <= pi/2

OpenStudy (anonymous):

woo.... quite tough..... So let's put aside this question for now. So the next question ask what value of a will h(x) have an inflection point at x=1....... So this would be something related to previous question right?

OpenStudy (amistre64):

h'' = 0, or undefined are critical points for inflection

OpenStudy (anonymous):

THen as you have said, C asks which values of a makes h(x) strictly increasing. So that would mean the things that we have been previously talking about right?

OpenStudy (amistre64):

that would tie in to A and B yes

OpenStudy (amistre64):

h' >= 0 is always increasing

OpenStudy (anonymous):

So for the inflection question, we need to find the value of a that makes h''=-a^2cos(ax) 0 right? then 0 would be the answer ?

OpenStudy (amistre64):

h'' = 0 for inflection @ x=1 yes

OpenStudy (amistre64):

brb

OpenStudy (amistre64):

recall that a is not 0 and not 4, but between 0 and 4, unless you misposted that information

OpenStudy (anonymous):

oh right... That means I would need to be finding another value that fits this...

OpenStudy (amistre64):

ill have to think about this and come back to it, most likely develop an idea tonight and post it tomorrow

OpenStudy (anonymous):

okay thanks I will be waiting

OpenStudy (anonymous):

@amistre64 Still no clue?

OpenStudy (amistre64):

nothing elementary :) i considered csc(a) = a, which is the intesection of y = csc(a) and y=a since y=a has a slope of 1, i thought; csc'(a) = 1 and came up with the golden ratio: x = (1 +- sqrt5)/2 so i thought maybe by finding that point i could interpolate for csc(a) = a

OpenStudy (amistre64):